OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 answers choices.. 2 -2 1 -1

OpenStudy (anonymous):

Which system of equations am I supposed to use. Which lesson is this assignment from?

OpenStudy (anonymous):

chapter 6, lesson 3.

OpenStudy (anonymous):

Which system of equations do I use? Substitution or Elimination!

OpenStudy (anonymous):

Ok. I'll use substitution and I'm working on the problem right now!

OpenStudy (anonymous):

2x + 3y=4 1st sol x - 2y = -5 2nd sol multiply 2nd sol by 2 and use elimination method

OpenStudy (anonymous):

So, Kantharaj! It's easier to use the elimination process not substitution?

OpenStudy (anonymous):

It's personal preference. I always do substitution.

OpenStudy (anonymous):

Yeah, me too. Thanks for that, ChmE!

OpenStudy (anonymous):

@alienababy How are you doing? did you get an answer?

OpenStudy (anonymous):

@ChmE, im working on it. Ive got y= 14, but im looking for x. I dont know what to do now?

OpenStudy (anonymous):

@alienbaby, you don't need my help anymore? I was almost finished with the equation!

OpenStudy (anonymous):

my bad, @alienababy

OpenStudy (anonymous):

@Johnjakile1998, i do! im trying to learn.

OpenStudy (anonymous):

hold on. Your y is wrong I got 2

OpenStudy (anonymous):

You don't have to shout! Lol. Alright, I'll put down the steps and then tell if you understand, and you can tell me.

OpenStudy (anonymous):

@Chme. I got the same y, @alienababy it's 2!

OpenStudy (anonymous):

im so confused. how?

OpenStudy (anonymous):

@ChmE, do you want to explain or me?

OpenStudy (anonymous):

@Johnjakile1998 will show the steps for the substitution and I'll show the steps for elimination

OpenStudy (anonymous):

sounds good to me :)

OpenStudy (anonymous):

Cool! I do better with substitution anyway! Quick question thought @ChmE. You substitute the first equation into the second one right

OpenStudy (anonymous):

I did 2nd into first cuz there is already an x alone so it made it quicker

OpenStudy (anonymous):

Yeah. Thanks, don't know how I missed it.

OpenStudy (anonymous):

\[2x+3y=4\]\[x-2y=-5\]We need to get one of the variables with the same coefficient. So lets multiply the second equation by 2.\[2x-4y=-10\]Now that we have a 2x in both equations we can ELIMINATE that term by subtracting the equations old school. One on top of the other and subtract like terms.\[7y=14\]2x-2x=0, 3x-(-4x)=7x, 4-(-10)=14 Now we solve for y and get 2. Plugging it into the 2nd equation because it is simpler we get -1 for x

OpenStudy (anonymous):

(3y**-(-4y**)=7y**)

OpenStudy (anonymous):

2x + 3y = 4 x - 2y = (-5) +2y +2y x = -5 + 2y 2(-5 + 2y) + 3y = 4 -10 + 4y + 3y = 4 +10 +10 4y + 3y = 14 7y = 14 -- --- 7 7 y = 2

OpenStudy (anonymous):

2x + 3y = 4 2x + 3(2) = 4 2x + 6 = 4 - 6 -6 2x = -2 -- --- 2 2 x = (-1)

OpenStudy (anonymous):

So, the solution is (-1,2) The value of x is (-1)

OpenStudy (anonymous):

@ChmE, so is my answer correct? I think so!

OpenStudy (anonymous):

How would you get x by itself if there is a number in front of it, like lets say.. 5x + 4y = 1 ?

OpenStudy (anonymous):

You can't x by itself, so you would have to do this:

OpenStudy (anonymous):

you can... x=(-4y+1)/5

OpenStudy (anonymous):

5x + 4y = 1 -5x -5x 4y = 1 - 5x And solve your problem using that equation!

OpenStudy (anonymous):

@ChmE, is that right?

OpenStudy (anonymous):

ya\[x=-\frac{ 4y }{ 5 }+\frac{ 1 }{ 5 }\]You are just solving for x even though there is another unknown. It is possible to get it by itself but not get a value. You always need the same amount of equations as unknowns to solve for a value.

OpenStudy (anonymous):

Thanks, @ChmE. @alienababy needs help on another problem. Are you going to help with it?

OpenStudy (anonymous):

@Johnjakile1998 & @ChmE ; What is the value of the y variable in the solution to the following system of equations? 5x+4y=1 4x+3y=-1

OpenStudy (anonymous):

Are you ready for this @ChmE?

OpenStudy (anonymous):

I'm solving the problem @alienababy

OpenStudy (anonymous):

thankyou @Johnjakile1998

OpenStudy (anonymous):

Welcome! While, I'm solving this problem. I wanted to know if you needed help on another question?

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x + y = 3 -3x + 4y = 1

OpenStudy (anonymous):

Which do I solve first?

OpenStudy (anonymous):

the first one i gave you.

OpenStudy (anonymous):

Cool. Come on @ChmE!

OpenStudy (anonymous):

For the above question for which I was mentioned in. I would use the elimination method because both variables have a coefficient. It will begin to get messy by substitution illustrated by me solving for x just a couple lines earlier. I'm going to let you attempt it first. If you follow my instructions for this method in the beginning I'm confident you'll get the right answer.

OpenStudy (anonymous):

@ChmE which one should i multiply first?

OpenStudy (anonymous):

Thanks! That's why I'm so lost.

OpenStudy (anonymous):

like in fractions find the lowest common multiple. In this case it is 12. The above equation needs to be multiplied by 3 and the bottom by 4. remember to multiple every term

OpenStudy (anonymous):

@alienababy. I'm going to solve the second equation first! I already got the answer!

OpenStudy (anonymous):

multiply**

OpenStudy (anonymous):

okay :) @Johnjakile1998 & @ChmE i got 15x+12y=3 then 16x+12=-4. is that right?

OpenStudy (anonymous):

For which problem?

OpenStudy (anonymous):

mine

OpenStudy (anonymous):

the first one, using the elimination method.

OpenStudy (anonymous):

its correct now subtract old school

OpenStudy (anonymous):

You're so weird @ChmE! Lol

OpenStudy (anonymous):

i got -x+y=-1 ? that doesnt seem right to me.

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

you shouldn't get a y term try again

OpenStudy (anonymous):

-x=-1 @ChmE

OpenStudy (anonymous):

@alienababy. I got the answer for the second equation. Want me to post it?

OpenStudy (anonymous):

yes please @Johnjakile1998

OpenStudy (anonymous):

what is -4-3?

OpenStudy (anonymous):

-7 @ChmE

OpenStudy (anonymous):

-7.

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

you can put whatever equation u want on top when subtracting. i put the 16x equation on top so that my x would be positive after subtraction.

OpenStudy (anonymous):

eliminates possible stupid mistakes

OpenStudy (anonymous):

oh.. @ChmE i didnt know that.

OpenStudy (anonymous):

if you want to see it for yourself. do the subtraction both ways. you should get -x=7 and x=-7. those or equal, no?

OpenStudy (anonymous):

yes, i did. @ChmE

OpenStudy (anonymous):

are**. Nice

OpenStudy (anonymous):

now just plug ur value of x into one of the equations, choosing the easier one (why make it even harder), and solve for y

OpenStudy (anonymous):

the first equations, or the second equations after you solved it? @ChmE

OpenStudy (anonymous):

whatever you want in this case. both equations look the same to me. It was more of a general statement

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

you want to use the original equations because it is simpler. smaller numbers

OpenStudy (anonymous):

2x + y = 3 -2x -2x y = 3 - 2x -3x + 4(3 - 2x) = 1 -3x + 12 - 8x = 1 -11x + 12 = 1 - 12 -12 -11x = -11 x = 1 2x + y = 3 2(1) + y = 3 2 + y = 3 -2 -2 y = 1 So, the solution is (1,1)

OpenStudy (anonymous):

gotcha. so it doesnt matter though? @ChmE

OpenStudy (anonymous):

no. You should get the same answer in either case. I'd encourage you to try it to see for yourself.

OpenStudy (anonymous):

thankyou @Johnjakile1998

OpenStudy (anonymous):

So, @alienababy that's the answer for the second equation. Is the answer right @ChmE?

OpenStudy (anonymous):

Any other questions @alienababy

OpenStudy (anonymous):

looks correct to me

OpenStudy (anonymous):

@ChmE i got -9?

OpenStudy (anonymous):

thanks @ChmE! You are @alienababy, just Algebra 1 isn't the math that is you're strongest. For all we know your strongest math is Calculus or Alebra 2 even Geometry!

OpenStudy (anonymous):

thanks :) @Johnjakile1998

OpenStudy (anonymous):

any others questions @alienababy

OpenStudy (anonymous):

Just wait.. Here's a problem I had on my last test\[y-\ln y=x ^{2}+1\]is this a solution of \[\frac{ dy }{ dx }=\frac{ 2xy }{ y-1 }?\]

OpenStudy (anonymous):

What kinda math is this?

OpenStudy (anonymous):

I got positive 9

OpenStudy (anonymous):

Differential Equations (Calc 4)